# Decentralization of Ethereum’s Builder Market

### v4

@@ Acknowledgements
This work is partially supported by an Ethereum Foundation Grant. We thank ultra sound relay for sharing the full bid dataset. We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements. We are grateful for the insightful comments by Prof. Campbell Harvey and the valuable discussions with Thomas Thiery, Barnabé Monnot, Danning Sui, Burak Öz, Data Always, and Georgios Konstantopoulos.
@@ Appendix E Private Order Flows
E.1 Pivotal Private Order Flow Providers
E.2 Requirements of Providers
@@ E.3 Minimal Cost for 1% Market Share
Appendix F Limitations of Our Approach and Data
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools [41, 42] and previous works [43, 44]. This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate.
Second, this approach may neglect off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study of evaluating competitiveness and efficiency, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
Third, this approach may overestimate builders’ profits because builders might refund providers for their private order flows. Such refunds may occur through private, off-chain deals and are therefore not captured. Overestimating the winner’s true value does not impact the competitiveness calculation; however, overestimating the second-highest true value could lead us to misclassify a competitive auction as uncompetitive. However, our study indicates that this kind of integration between top builders and providers began on a large scale starting in 2024, after the time range of our evaluation of competitiveness and efficiency.
Missing pivotal providers.   Our results in Sec. 5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies [18, 19]. However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of providers will only be underestimated.
Our auction dataset might miss partial bids since they are self-reported by relays. Recall that we used the second-highest bid value in an MEV-Boost auction as the threshold to determine a pivotal provider. Therefore, underestimating the second-highest bid value might also lead us to underestimate the significance of each provider.

### v5

@@ Acknowledgements
This work is partially supported by an Ethereum Foundation Grant. We thank our shepherd and the reviewers for their helpful comments, which significantly improved the paper. We are grateful to ultra sound relay for sharing the full bid dataset. We also thank the teams at BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions regarding access requirements. Finally, we appreciate the insightful comments from Prof. Campbell Harvey and the valuable discussions with Thomas Thiery, Barnabé Monnot, Danning Sui, Burak Öz, Data Always, and Georgios Konstantopoulos.
@@ Appendix E Private Order Flows
E.1 Pivotal Private Order Flow Providers
E.2 Requirements of Providers